OpenStudy (anonymous):

A ski slope has an incline of 64 (its angle of elevation is 64). The slope has an elevation of 3873 feet. At the bottom of the slope is a flat (horizontal) run out, and 898 feet from the slope is the end of a cable (the other end is at the top of the slope.) How long is the cable?

OpenStudy (anonymous):

as we compute sin 64= 0.899 and cos 64=0.438 consider triangle ABC ,AB =AC*sin 64=3873*0.899=3481.27 and BC=AC*cos 64=3873*0.438=1696.374 DB=DC+CB=898+1696.374 so tan D=AB/DB=3481.27/2594.374=1.314 so D=52.728 degrees( using calculator) so AB=AD*sin52.728 or AD=AB/sin 52.728 AD= 4374.72 m is the cable length